R is the region bounded by x = y


2 + 1 and x = 9 − y


2


and the axis


of revolution is x = 9.

1
Expert's answer
2021-11-01T18:45:38-0400




"y^2+1=9-y^2"

"2y^2=8"

"y^2=4"

"y_1=-2, y_2=2"

"A(x)=\pi(outer\ radius)^2-\pi (inner\ radius)^2"

"=\pi (9-(y^2+1))^2-\pi(9-(9-y^2))^2"

"=\pi(81-18y^2-18+y^4+2y^2+1"

"-81+162-18y^2-81+18y^2-y^4)"

"=\pi(-16y^2+64)"


"V=\displaystyle\int_{-2}^2\pi(-16y^2+64)dy=32\pi\displaystyle\int_{0}^2\pi(-y^2+4)d"

"=32\pi\bigg[-\dfrac{y^3}{3}+4y\bigg]\begin{matrix}
 2 \\
 0
\end{matrix}=32\pi(-\dfrac{8^3}{3}+8)"

"=\dfrac{512\pi}{3} (units^3)"